# R-matrix valued Lax pair for elliptic Calogero-Inozemtsev system

### v2

@@ Abstract
We consider the elliptic Calogero-Inozemtsev system of ${\rm BC}_{n}$ type with five arbitrary constants and propose $R$-matrix valued generalization for $2n\times 2n$ Takasaki’s Lax pair. For this purpose we extend the Kirillov’s ${\rm B}$-type associative Yang-Baxter equations to the similar relations depending on the spectral parameters and the Planck constants. General construction uses the elliptic Shibukawa-Ueno $R$-operator and the Komori-Hikami $K$-operators satisfying reflection equation. Then, using the Felder-Pasquier construction the answer for the Lax pair is also written in terms of the Baxter’s 8-vertex $R$-matrix. As a by-product of the constructed Lax pair we also propose ${\rm BC}_{n}$ type generalization for the elliptic XYZ long-range spin chain, and we present arguments pointing to its integrability.
@@ 1 Introduction
The elliptic Calogero-Moser model of ${\rm gl}_{n}$ type is an integrable systems of classical mechanics, describing pairwise interaction of $n$ particles. It is defined by the Hamiltonian function [25]:
@@
where $\wp(z)$ is the Weierstrass Weierstrass-$\wp$-function, $g\in\mathbb{C}$ is a coupling constant, $p_{i}\in\mathbb{C}$ and $q_{i}\in\mathbb{C}$ are the particles momenta and positions. The Poisson brackets are canonical:
@@
$\begin{array}[]{c}\displaystyle{{\dot{q}}_{i}=p_{i}\,,\quad{\ddot{q}}_{i}={g}^{2}\sum\limits_{k:k\neq i}^{n}\wp^{\prime}(q_{ik})}\end{array}$ (1.3)
@@
This relation is a particular case of a more general one known as the associative Yang-Baxter equation (AYBE) [7, 27]:
$\begin{array}[]{c}\displaystyle{R^{z}_{12}R^{w}_{23}=R^{w}_{13}R_{12}^{z-w}+R^{w-z}_{23}R^{z}_{13}\,,\ \ R^{z}_{ij}=R^{z}_{ij}(q_{i},q_{j})\,,}\end{array}$ (1.8)
where $R^{z}_{12}(x_{1},x_{2})$ is an operator that acts on a product of vector spaces. For $R$-matrices depending on the difference of spectral parameters $R^{z}_{12}(x_{1},x_{2})=R^{z}_{12}(x_{1}-x_{2})$ and in the case where the action on vector spaces is by multiplication on the function $\phi(z,x_{1}-x_{2})$ we reproduce (1.7). In the general case, equation (1.8) is fulfilled by the elliptic Baxter-Belavin $R$-matrix [27].
It can be show (see e.g. [21]) that any solution of (1.8) satisfying certain additional properties (see below the unitarity (2.5) the and skew-symmetry (2.6)) satisfies also the quantum Yang-Baxter equation
$\begin{array}[]{l}\displaystyle{R^{\hbar}_{12}(x_{1},x_{2})R^{\hbar}_{13}(x_{1},x_{3})R^{\hbar}_{23}(x_{2},x_{3})=R^{\hbar}_{23}(x_{2},x_{3})R^{\hbar}_{13}(x_{1},x_{3})R^{\hbar}_{12}(x_{1},x_{2})\,.}\end{array}$ (1.9)
@@
$\begin{array}[]{c}\displaystyle{{\mathcal{L}}_{ij}(z)=\delta_{ij}p_{i}+{g}(1-\delta_{ij})R^{z}_{ij}(q_{i}-q_{j})\,.}\end{array}$ (1.10)
@@
where $\omega_{\gamma}$ are half-periods (A.18), and the five arbitrary constants are $g,\nu_{0},\nu_{1},\nu_{2},\nu_{3}\in\mathbb{C}$. Originally, the Lax representation was of size $3n\times 3n$. We use $2n\times 2n$ Lax representation proposed by K. Takasaki in [31]. More precisely, we use it in the form obtained by O. Chalykh in [3] as a limiting case of a more general relativistic (Ruijsenaars-van Diejen) model.
In our construction of the $R$-matrix valued Lax pair we also use the $K$-matrices [12], which are solutions to the reflection equation [30]:
$\begin{array}[]{c}\displaystyle{R^{-}_{12}(x_{1},x_{2})K^{\hbar}_{1}(x_{1})R^{+}_{12}(x_{1},x_{2})K^{\hbar}_{2}(x_{2})=K^{\hbar}_{2}(x_{2})R^{+}_{12}(x_{1},x_{2})K^{\hbar}_{1}(x_{1})R^{-}_{12}(x_{1},x_{2})\,,}\end{array}$ (1.12)
where $R^{-}_{12}(x_{1},x_{2})=R^{\hbar}_{12}(x_{1}-x_{2})$, $R^{+}_{12}(x_{1},x_{2})=R^{\hbar}_{12}(x_{1}+x_{2})$. The $K$-matrices play the role of the boundary conditions in quantum integrable systems.
In [11, 16, 17] Y. Komori and K. Hikami suggested solution of (1.12) with the operator valued Shibukawa-Ueno $R$-matrix in the form of certain operators ($K$-operators). This was applied for constructing commuting set of quantum Hamiltonians for the model (1.11) and its relativistic generalization (through the Dunkl-Cherednik type operators). We actively use the results of [11, 16, 17] in our consideration.
@@
Then (1.8) can be considered as representation of (1.13) with parameters. In [18] the set of relations was extended to other root systems. For the root system of type $B$ the set of relations includes the four term relation
@@
It was explained by G. Felder and V. Pasquier in [6] that the Shibukawa-Ueno $R$-operator can be transformed into the elliptic Baxter-Belavin $R$-matrix by specifying the space, where this operator acts. In a similar way the $K$-operators from [11, 16] were transformed into the matrix form in [17]. We use these results and prove that the obtained relations for the ${\rm BC}_{n}$ associative Yang-Baxter equation with parameters are valid for the Baxter’s 8-vertex $R$-matrix and the elliptic $K$-matrix.
Finally, using the approach of [28] we present the quantum Hamiltonian for anisotropic long-range spin chain of ${\rm BC}_{n}$ type.
@@ 2.1 Krichever’s Lax pair
Consider the Lax pair for the elliptic ${\rm gl}_{n}$ Calogero-Moser model (1.1)-(1.3). The Lax pair of size $n\times n$, and the Lax matrix has the form (1.5) Similarly, for the matrix elements of $M$-matrix:
@@
See Appendix for elliptic functions definitions and properties. The proof of the Lax equation (1.4) with the Lax pair (1.5), (2.1) is based on the identities (A.15)-(A.17) written as
@@
These relations follow from the (genus one Fay identity) summation formula (A.12):
@@ 2.2 $R$-matrix identities
$\begin{array}[]{l}\displaystyle{R_{12}^{\hbar}(z)R_{21}^{\hbar}(-z)=(\wp(\hbar)-\wp(z))\,.}\end{array}$ (2.5)
This is an $R$-matrix analogue of (A.16). Also, we have a skew-symmetry property
${R^{\hbar}_{12}(z)=-R^{-\hbar}_{21}(-z)\,,}$ (2.6)
which is an analogue of the simple relation $\phi(\hbar,z)=-\phi(-\hbar,-z)$. Main relation is the associative Yang-Baxter equation [7]
@@
$\begin{array}[]{c}\displaystyle{R^{z}_{ij}F^{z}_{ji}-F^{z}_{ij}R^{z}_{ji}=\wp^{\prime}(q_{ij})\,.}\end{array}$ (2.11)
@@ 2.3 $R$-matrix valued Lax pairs
$\begin{array}[]{c}\displaystyle{{{\mathcal{L}}}(z)=\sum\limits_{i,j=1}^{n}E_{ij}\otimes{\mathcal{L}}_{ij}(z)\,,\quad\quad{\mathcal{L}}_{ij}(z)=1_{N}^{\otimes n}\,\delta_{ij}p_{i}+g(1-\delta_{ij})R^{\,z}_{ij}(q_{ij})}\end{array}$ (2.12)
¯$\begin{array}[]{c}\displaystyle{\bar{\mathcal{M}}_{ij}(z)=D_{i}\delta_{ij}+g(1-\delta_{ij})F^{\,z}_{ij}(q_{ij})+\delta_{ij}\,{\mathcal{F}}^{\,0}\,,\quad D_{i}=-g\sum\limits_{k:k\neq i}^{n}F^{\,0}_{ik}(q_{ik})\,,}\end{array}$ (2.13)
@@
[…]
The proof is based on the identities (2.10)-(2.11) and the following simple [^2] Relation (2.15) does not use any specific $R$-matrix properties but only $[R_{ab}^{z},F^{w}_{cd}]=0$ for all distinct $a,b,c,d$. relation:
@@
Similarly, $R$-matrix valued Lax pairs can be deduced for other root systems. Such results were described in [10] by extending the Lax pairs [25, 4] for the Calogero-Moser models related to the classical root systems of $B,C,D$ types. In accordance with the original construction from [25, 4] in these cases we have not arbitrary coupling constants, they satisfy a set of constraints. In [22] the ${BC}_{1}$ case was considered with four arbitrary constants. In both papers [10] and [22] an additional component of the quantum Hilbert space was used in order to define matrix elements ${\mathcal{L}}_{i,i+n}$ and ${\mathcal{L}}_{n+i,i}$. Below we use different approach based on the usage of $K$-matrices.
@@ 2.4 Long-range spin chains
Having an $R$-matrix valued Lax pair one can deduce a quantum long-range spin chain in the following way [28]. Denote by ${{\mathcal{M}}}(z)$ the part of ¯$\bar{\mathcal{M}}(z)$ without the term ${\mathcal{F}}^{0}$, that is
¯$\begin{array}[]{c}\displaystyle{\bar{\mathcal{M}}(z)={\mathcal{M}}(z)+1_{n}{\mathcal{F}}^{0}\,.}\end{array}$ (2.16)
@@
˙$\begin{array}[]{c}\displaystyle{{\dot{\mathcal{L}}}(z)=[{\mathcal{L}}(z),\bar{\mathcal{M}}(z)]}\end{array}$ (2.17)
@@
$\begin{array}[]{c}\displaystyle{q_{j}\rightarrow x_{j}=\frac{j}{n}\,,\quad j=1,...,n\,.}\end{array}$ (2.18)
@@
$\begin{array}[]{c}\displaystyle{{\mathcal{L}}^{\prime}(z)={\mathcal{L}}(z)|_{q_{j}=x_{j},p_{j}=0}\,,\quad{\mathcal{M}}^{\prime}(z)={\mathcal{M}}(z)|_{q_{j}=x_{j}}\,,\quad{{\mathcal{F}}^{0}}^{\prime}={\mathcal{F}}^{0}|_{q_{j}=x_{j}}\,.}\end{array}$ (2.19)
Then from the Lax equation (taking into account that ˙${\dot{\mathcal{L}}}(z)|_{q_{j}=x_{j}}=0$) we get
@@
This construction can be regarded as an explanation of the Polychronakos freezing trick, which relates the quantum spin Calogero-Moser models with the long-range spin chains by restricting to the classical equilibrium position of the corresponding spinless model. The above idea was used in [28] to obtain anisotropic version of the Inozemtsev long-range chain [14]. Let us also mention that isotropic trigonometric ${\rm BC}_{n}$ type long-range spin chains were described in [5] using the Dunkl operator formalism.
Different generalizations and applications for this type of models can be found in [20, 24, 15] and references therein.

### v3

@@ Abstract
We consider the elliptic Calogero-Inozemtsev system of ${\rm BC}_{n}$ type with five arbitrary constants and propose $R$-matrix valued generalization for $2n\times 2n$ Takasaki’s Lax pair. For this purpose, we extend the Kirillov’s $\rm B$}-type associative Yang-Baxter equations to similar relations depending on the spectral parameters and the Planck constants. General construction uses the elliptic Shibukawa-Ueno $R$-operator and the Komori-Hikami $K$-operators satisfying the reflection equation. Then, using the Felder-Pasquier construction, the answer for the Lax pair is also written in terms of the Baxter’s 8-vertex $R$-matrix. As a by-product of the constructed Lax pair we also propose a ${\rm BC}_{n}$ type generalization for the elliptic XYZ long-range spin chain, and we present arguments pointing to its integrability.
@@ 1 Introduction
The elliptic Calogero-Moser model of ${\rm gl}_{n}$ type is an integrable system of classical mechanics, describing pairwise interaction of $n$ particles. It is defined by the Hamiltonian function [26]:
@@
where $\wp(z)$ is the Weierstrass $\wp$-function, $g\in\mathbb{C}$ is a coupling constant, $p_{i}\in\mathbb{C}$ and $q_{i}\in\mathbb{C}$ are the particles momenta and positions. The Poisson brackets are canonical:
@@
$\begin{array}[]{c}\displaystyle{{\dot{q}}_{i}=p_{i}\,,\quad{\ddot{q}}_{i}={g}^{2}\sum\limits_{k:k\neq i}^{n}\wp^{\prime}(q_{i}-q_{k})}\end{array}$ (1.3)
@@
This relation is a particular case of a more general one known as the associative Yang-Baxter equation (AYBE) introduced by S. Fomin and An. Kirillov in [7]. Let $\mathcal{A}$} be an associative algebra and $R^{z}(q_{1},q_{2})$ be a meromorphic function depending on parameters $z,q_{1},q_{2}\in\mathbb{C}$ with values in $\mathcal{A}\otimes\mathcal{A}$. The AYBE with parameters proposed in [28] is the following equation in $\mathcal{A}\otimes\mathcal{A}\otimes\mathcal{A}$:
$\begin{array}[]{c}\displaystyle{R^{z}_{12}(q_{1},q_{2})R^{w}_{23}(q_{2},q_{3})=R^{w}_{13}(q_{1},q_{3})R_{12}^{z-w}(q_{1},q_{2})+R^{w-z}_{23}(q_{2},q_{3})R^{z}_{13}(q_{1},q_{3})\,,}\end{array}$ (1.8)
where $R^{z}_{ij}(q_{i},q_{j})$ means the operator acting nontrivial as $R^{z}(q_{i},q_{j})$ on the $i$-th and the $j$-th tensor components. In case when $R^{z}(q_{1},q_{2})$ depends on the difference of spectral parameters $R^{z}(q_{1},q_{2})=R^{z}(q_{1}-q_{2})$ and $\mathcal{A}=\mathbb{C}$ the AYBE (1.8) reproduces (1.7), where $R^{z}(q_{1},q_{2})$ acts by multiplication by the function $\phi(z,q_{1}-q_{2})$. In the matrix case $\mathcal{A}={\rm Mat}(N,\mathbb{C})$, the equation (1.8) is fulfilled by the elliptic Baxter-Belavin $R$-matrix [28].
It can be shown (see e.g. [21]) that any solution of (1.8) satisfying certain additional properties (see below the unitarity (2.5) the and skew-symmetry (2.6)) satisfies also the quantum Yang-Baxter equation
$\begin{array}[]{l}\displaystyle{R^{\hbar}_{12}(q_{1},q_{2})R^{\hbar}_{13}(q_{1},q_{3})R^{\hbar}_{23}(q_{2},q_{3})=R^{\hbar}_{23}(q_{2},q_{3})R^{\hbar}_{13}(q_{1},q_{3})R^{\hbar}_{12}(q_{1},q_{2})\,.}\end{array}$ (1.9)
@@
$\begin{array}[]{c}\displaystyle{{\mathcal{L}}_{ij}(z)={\rm Id}\,\delta_{ij}p_{i}+{g}(1-\delta_{ij})R^{z}_{ij}(q_{i}-q_{j})\in{\rm Mat}(N,\mathbb{C})^{\otimes n}\,.}\end{array}$ (1.10)
@@
where $\omega_{\gamma}$ are half-periods (A.18), and the five arbitrary constants are $g,\nu_{0},\nu_{1},\nu_{2},\nu_{3}\in\mathbb{C}$. Originally, the Lax representation was of size $3n\times 3n$. We use $2n\times 2n$ Lax representation proposed by K. Takasaki in [33]. More precisely, we use the Lax pair in the form obtained by O.Chalykh in a conceptual way from Cherednik-Dunkl operators in [3].
In our construction of the $R$-matrix valued Lax pair we also use $K$-matrices $K^{\hbar}(q)\in{\rm Mat}(N,\mathbb{C})$ [12] , which are solutions to the reflection equation [32]:
$\begin{array}[]{c}\displaystyle{R^{-}_{12}(q_{1},q_{2})K^{\hbar}_{1}(q_{1})R^{+}_{12}(q_{1},q_{2})K^{\hbar}_{2}(q_{2})=K^{\hbar}_{2}(q_{2})R^{+}_{12}(q_{1},q_{2})K^{\hbar}_{1}(q_{1})R^{-}_{12}(q_{1},q_{2})\,,}\end{array}$ (1.12)
where $R^{-}_{12}(q_{1},q_{2})=R^{\hbar}_{12}(q_{1}-q_{2})$, $R^{+}_{12}(q_{1},q_{2})=R^{\hbar}_{12}(q_{1}+q_{2})$. The $K$-matrices play the role of the boundary conditions in quantum integrable systems.
In [11, 16, 17] Y. Komori and K. Hikami suggested a solution of (1.12) with the operator valued Shibukawa-Ueno $R$-matrix in the form of certain operators ($K$-operators). This was applied to construct a commuting set of quantum Hamiltonians for the model (1.11) and its relativistic generalization (through the Dunkl-Cherednik type operators). We actively use the results of [11, 16, 17] in our consideration.
@@
Then (1.8) can be considered as a representation of (1.13) with parameters. In [18] the set of relations was extended to other root systems. For the root system of type $B$ the set of relations includes the four term relation
@@
G. Felder and V. Pasquier in [6] explained that the Shibukawa-Ueno $R$-operator can be transformed into the elliptic Baxter-Belavin $R$-matrix by specifying the space, where this operator acts. Similarly, the $K$-operators from [11, 16] were transformed into the matrix form in [17]. We use these results and prove that the obtained relations for the ${\rm BC}_{n}$ associative Yang-Baxter equation with parameters are valid for the Baxter’s 8-vertex $R$-matrix and the elliptic $K$-matrix.
Finally, using the approach of [30] we present the quantum Hamiltonian for an anisotropic long-range spin chain of ${\rm BC}_{n}$ type.
A list of notation used in the paper is given before the Appendix.
@@ 2 Elliptic Calogero-Moser model and $R$-matrices
In this Section we recall main steps of construction of long-range spin chains from [30] based on the $R$-matrix valued Lax pairs [21]. We begin with the famous classical Krichever’s Lax pair [19]. Then we describe the necessary properties of $R$-matrices which are then used as matrix elements in the $R$-matrix valued Lax pairs. Finally, we explain that at equilibrium positions, the classical Lax equation can be treated as a quantum Lax equation for some long-range spin chain.
@@ 2.1 Krichever’s Lax pair
Consider the Lax pair for the elliptic ${\rm gl}_{n}$ Calogero-Moser model (1.1)-(1.3). The Lax pair of size $n\times n$, and the Lax matrix has the form (1.5). Similarly, for the matrix elements of $M$-matrix:
@@
where $q_{ij}$ means $q_{i}-q_{j}$. See the Appendix for elliptic functions definitions and properties. The proof of the Lax equation (1.4) with the Lax pair (1.5), (2.1) is based on the identities (A.15)-(A.17) written as
@@
These relations follow from the summation formula (genus one Fay identity) (A.12):
@@ 2.2 $R$-matrix identities
Let us recall the main properties of the elliptic $R$-matrix which will be used for the construction of the $R$-matrix valued Lax pair. In what follows, we deal with the Shibukawa-Ueno (operator-valued) and the Baxter-Belavin (matrix-valued) $R$-matrices. These properties are valid for both versions.
@@
$\begin{array}[]{l}\displaystyle{R_{12}^{\hbar}(u)R_{21}^{\hbar}(-u)=(\wp(\hbar)-\wp(u)){\rm Id}\,,}\end{array}$ (2.5)
where $\rm Id$} is either an identity operator or an identity matrix. This is an $R$-matrix analogue of (A.16). Also, we have a skew-symmetry property
${R^{\hbar}_{12}(u)=-R^{-\hbar}_{21}(-u)\,,}$ (2.6)
which is an analogue of the simple relation $\phi(\hbar,u)=-\phi(-\hbar,-u)$. The main relation is the associative Yang-Baxter equation [7]
@@
$\begin{array}[]{c}\displaystyle{R^{z}_{ij}F^{z}_{ji}-F^{z}_{ij}R^{z}_{ji}=\wp^{\prime}(q_{ij}){\rm Id}\,.}\end{array}$ (2.11)
@@ 2.3 $R$-matrix valued Lax pairs
$\begin{array}[]{c}\displaystyle{{{\mathcal{L}}}(z)=\sum\limits_{i,j=1}^{n}E_{ij}\otimes{\mathcal{L}}_{ij}(z)\,,\quad\quad{\mathcal{L}}_{ij}(z)={\rm Id}\,\delta_{ij}p_{i}+g(1-\delta_{ij})R^{\,z}_{ij}(q_{ij})}\end{array}$ (2.12)
$\begin{array}[]{c}\displaystyle{\bar{\mathcal{M}}_{ij}(z)=D_{i}\delta_{ij}+g(1-\delta_{ij})F^{\,z}_{ij}(q_{ij})+\delta_{ij}\,{\mathcal{F}}^{\,0}\,,\quad D_{i}=-g\sum\limits_{k:k\neq i}^{n}F^{\,0}_{ik}(q_{ik})\,,}\end{array}$ (2.13)
@@
 $\begin{array}[]{c}\displaystyle{{\mathcal{F}}^{\,0}=g\sum\limits_{k>m}^{n}F^{\,0}_{km}(q_{km})=\frac{g}{2}\sum\limits_{k\neq m}^{n}F^{\,0}_{km}(q_{km})\,.}\end{array}$ (2.14)
If an operator $R^{z}(q)$ satisfies the properties from the previous section, namely (2.5), (2.6), (2.7), the classical Lax equation ${\dot{\mathcal{L}}}(z)=[{\mathcal{L}}(z),\bar{\mathcal{M}}(z)]$ is equivalent to the classical equations of motion (1.3) for the Calogero-Moser system. The proof of this fact is based on the identities (2.10)-(2.11) and the following simple [^2] Relation (2.15) does not use any specific $R$-matrix properties but only $R_{ab}^{z},F^{w}_{cd}]=0$ for all distinct $a,b,c,d$. relation:
@@
As a result, one gets the classical equations of motion (1.3). That is (2.12)-(2.13) is just a multi-dimensional generalization of the Krichever’s Lax pair (1.10), (2.1).
Originally, the $R$-matrix valued Lax pair was introduced using the elliptic Baxter-Belavin $R$-matrix for the ${\rm GL}_{N}$ Lie group, so that in (2.12)-(2.14) $R^{z}_{ij}(q_{ij})\in{\rm Mat}(N,\mathbb{C})^{\otimes n}$, $\rm Id$} is an identity matrix in ${\rm Mat}(N,\mathbb{C})^{\otimes n}$ and ${\mathcal{L}}(z)\in{\rm Mat}(n,\mathbb{C})\otimes{\rm Mat}(N,\mathbb{C})^{\otimes n}$. Notice that all constructions are valid for the Shibukawa-Ueno $R$-operator, since it satisfies all the properties from the previous section.
Similarly, $R$-matrix valued Lax pairs can be deduced for other root systems. Such results were described in [10] by extending the Lax pairs [26, 4] for the Calogero-Moser models related to the classical root systems of $B,C,D$ types. In accordance with the original construction from [26, 4] in these cases we have not arbitrary coupling constants, they satisfy a set of constraints. In [22] the ${BC}_{1}$ case was considered with four arbitrary constants. In both papers [10] and [22] an additional component of the quantum Hilbert space was used in order to define matrix elements ${\mathcal{L}}_{i,i+n}$ and ${\mathcal{L}}_{n+i,i}$. In the following, we use a different approach which is based on the usage of $K$-matrices.
@@ 2.4 Long-range spin chains
Here we assume that we deal with a matrix-valued $R$-operator: $R^{z}_{ij}(q_{ij})\in{\rm Mat}(N,\mathbb{C})^{\otimes n}$. Having a $R$-matrix valued Lax pair one can deduce a quantum long-range spin chain in the following way [30]. Denote by ${{\mathcal{M}}}(z)$ the part of $\bar{\mathcal{M}}(z)$ without the term ${\mathcal{F}}^{0}\in{\rm Mat}(N,\mathbb{C})^{\otimes n}$, that is
$\begin{array}[]{c}\displaystyle{\bar{\mathcal{M}}(z)={\mathcal{M}}(z)+1_{n}\otimes{\mathcal{F}}^{0}\,.}\end{array}$ (2.16)
@@
$\begin{array}[]{c}\displaystyle{{\dot{\mathcal{L}}}(z)=[{\mathcal{L}}(z),\bar{\mathcal{M}}(z)]}\end{array}$ (2.17)
@@
$\begin{array}[]{c}\displaystyle{q_{j}\rightarrow\zeta_{j}=\frac{j}{n}\,,\quad j=1,...,n\,.}\end{array}$ (2.18)
@@
$\begin{array}[]{c}\displaystyle{{\mathcal{L}}^{\prime}(z)={\mathcal{L}}(z)|_{q_{j}=\zeta_{j},p_{j}=0}\,,\quad{\mathcal{M}}^{\prime}(z)={\mathcal{M}}(z)|_{q_{j}=\zeta_{j}}\,,\quad{{\mathcal{F}}^{0}}^{\prime}={\mathcal{F}}^{0}|_{q_{j}=\zeta_{j}}\,.}\end{array}$ (2.19)
Then from the Lax equation (taking into account that ${\dot{\mathcal{L}}}(z)|_{q_{j}=\zeta_{j}}=0$) we get
@@
This construction can be regarded as an explanation of the Polychronakos freezing trick [29], which relates the quantum spin Calogero-Moser models with the long-range spin chains by restricting to the classical equilibrium position of the corresponding spinless model. The above idea was used in [30] to obtain anisotropic version of the Inozemtsev long-range chain [14]. Recently, an explanation of the freezing trick was suggested within the framework of hybrid integrable systems in [24]. Let us also mention that isotropic trigonometric ${\rm BC}_{n}$ type long-range spin chains were described in [5] using the Dunkl operator formalism.
Different generalizations and applications for this type of model can be found in [20, 25, 15] and references therein.